### v1

@@ Abstract
Distributions in spatial model often exhibit localized features. Intuitively, this locality implies a low intrinsic dimensionality, which can be exploited for efficient approximation and computation of complex distributions. However, existing approximation theory mainly considers the joint distributions, which does not guarantee that the marginal errors are small. In this work, we establish a dimension independent error bound for the marginals of approximate distributions. This $\ell_{\infty}$-approximation error is obtained using Stein’s method, and we propose a $\delta$-locality condition that quantifies the degree of localization in a distribution. We also show how $\delta$-locality can be derived from different conditions that characterize the distribution’s locality. Our $\ell_{\infty}$ bound motivates the localization of existing approximation methods to respect the locality. As examples, we show how to use localized likelihood-informed subspace method and localized score matching, which not only avoid dimension dependence in the approximation error, but also significantly reduce the computational cost due to the local and parallel implementation based on the localized structure.
@@ 1 Introduction
Approximation of distributions is a fundamental problem in applied mathematics and computer science. With the development of modern machine learning methods [15, 9, 21, 11], significant advancements are achieved in approximation of complex distributions. Compared to vast empirical success, concrete mathematical understanding of these methods is still lacking, especially in the context of high-dimensionality.
In this paper, we develop a $\ell_{\infty}$-approximation theory for a large class of high-dimensional distributions that exhibit localized features, explaining how such distributions can be approximated efficiently without suffering from the curse of dimensionality. These localized distributions are common in spatial models and graphical models arising from many statistical physics and computer science problems. A typical example is the Ising model where there are only local interactions.
Intuitively, such localized structure implies that the distribution is intrinsically low-dimensional, which can be exploited for efficient approximation and computation. We formalize this intuition and proves a dimension independent error bound for the marginals of approximate distributions. More precisely, we show in 2.4 that under some locality assumption, the $1$-Wasserstein marginal approximation error is bounded by local perturbations of the log density:
@@
$\max_{i}W_{1}(\pi_{i},\pi_{i}^{\prime})\leq\delta\cdot\max_{i}\int\left|\nabla_{i}\log\pi(x)-\nabla_{i}\log\pi^{\prime}(x)\right|\pi(x)\mathrm{d}x.$
Note that the choice of the log density is motivated by its foundational role in many state-of-the-art approximation and sampling methods, such as the score matching [14], variational autoencoder [15] and diffusion models [21, 11]. Approximation error in terms of the log density is well documented in the literature, but most only consider the joint distribution, which in general is dimension dependent. One typical example is the Otto-Villani inequality [19]:
@@
$W_{2}^{2}(\pi,\pi^{\prime})\leq c^{-2}\int\left|\nabla\log\pi(x)-\nabla\log\pi^{\prime}(x)\right|^{2}\pi(x)\mathrm{d}x.$
These bounds does not guarantee that the marginal error is small, and becomes less informative when one is only interested in the statistics of the marginals. Our $\ell_{\infty}$-approximation result fills the gap by exploiting the locality of the distribution and avoid the dimension dependence here.
Locality in the distributions is well-known and exploited in the communities like spatial statistics [25, 6, 20], data assimilation [13, 10] and image processing [27]. The key observation is that the correlation of parameters at large distances is neglectable in many spatial models. Therefore, modeling and computing such distributions can be done in an approximate manner, by artificially removing long-range correlations, thus greatly improving the computational efficiency. Such technique is called Vecchia approximation [25] in spatial statistics, and localization [13] in data assimilation. Despite the extensive application studies, the mathematical characterization of this approximation is rarely explored. With this in mind, we discuss in Section 2.3 the problem of approximating localized distributions by undirected graphical models (also known as the Markov random fields [16, 27]), i.e. one aims to find local potentials $u_{i}$ that depends only the neighboring sites $N_{i}$ of the site $i$, such that
$\pi(x)\approx\frac{1}{Z}\exp\left(\sum_{i}u_{i}(x_{N_{i}})\right).$
Note this is the prototype math formulation of the Vecchia approximation.
In the sampling literature, there are also recent work [17, 23, 8, 2] exploiting the locality for efficient sampling. [17] proposed to apply the localization technique in MCMC sampling to solve Bayesian inverse problems and obtained a dimension independent convergence rate. [23] extended this idea and proved a dimension independent convergence rate of the Metropolis-Adjusted-Langevin-Algorithm-within-Gibbs (MLwG) for distributions with a sparse condition structure (equivalent to a sparse graphical model). [8] made a initial trial to obtain $\ell_{\infty}$-approximation error for a image deblurring problem, based on which we further develop the theory to the general case in this paper. We also notice a concurrent work [2] that shows the marginals of unadjusted Langevin algorithm converges to the target at a convergence rate with logarithm dependence on the dimension. They introduced a $\ell_{\infty}$-type Wasserstein distance, which follows a similar focus on the marginals as in our work.
In this paper, we propose the concept of $\delta$-locality in 2.1 to quantifies how a distribution is localized. We will use the $\delta$-locality to obtain $\ell_{\infty}$-approximation error by applying the Stein’s method [22]. We show how to obtain $\delta$-locality given different characterization of the localized distribution, including the sparse graphical model. The main technical contribution is to show that $\delta$ is dimension independent for typical localized distributions. The $\ell_{\infty}$-approximation bound motivates to localize existing approximation methods to respect the locality. As examples, we show how to localize the likelihood-informed subspace (LIS) method [4, 3, 5] and the score matching method [14, 21, 11]. Such localized approximation method not only yields a dimension independent $\ell_{\infty}$ error bound, but also greatly reduces the computational cost due to the local and parallel implementation based on the localized structure.
The paper is organized as follows. The main results are presented in Section 2, where we define the $\delta$-locality, proves the main approximation theory, and then discuss some useful extensions. The following three sections discuss three applications of the theory: the localized LIS method in Section 3, the localized score matching in Section 4, and the localized neural network surrogate in Section 5. In Section 6, we provide the main steps of the technical results for the $\delta$-locality.
@@
  • Norms. For $v\in\mathbb{R}^{d}$, briefly denote the vector $2$-norm as $\left|v\right|$. Likewisely, for a matrix $A\in\mathbb{R}^{m\times n}$, $\left\|A\right\|=\sup_{v\neq 0}\frac{\left|Av\right|}{\left|v\right|}$ denotes the matrix $2$-norm. $\left\|A\right\|_{\rm F}$ denotes the Frobenius norm $\left\|A\right\|_{\rm F}^{2}=\sum_{i,j}A_{ij}^{2}$.
    For a function $f:\mathbb{R}^{d}\to\mathbb{R}$, denote $\left\|f\right\|_{L^{p}(\pi)}$ as the weighted $L^{p}$ norm $\left\|f\right\|_{L^{p}(\pi)}^{p}=\int_{\mathbb{R}^{d}}|f(x)|^{p}\pi(x)\mathrm{d}x$. For a vector-valued function $f:\mathbb{R}^{d}\to\mathbb{R}^{m}$, still denote $\left\|f\right\|_{L^{p}(\pi)}^{p}=\int_{\mathbb{R}^{d}}\left|f(x)\right|^{p}\pi(x)\mathrm{d}x$ and note here $\left|f(x)\right|$ is the vector $2$-norm without specification.
@@
  • Probability. Denote $\mathcal{P}_{1}(\mathbb{R}^{d})$ as the set of probability distributions in $\mathbb{R}^{d}$ with finite first moment. The $1$-Wasserstein distance between two distributions $\pi,\pi^{\prime}\in\mathcal{P}_{1}(\mathbb{R}^{d})$ is defined as
    $W_{1}(\pi,\pi^{\prime})=\inf_{\gamma\in\Pi(\pi,\pi^{\prime})}\int_{\mathbb{R}^{d}}\left|x-y\right|\mathrm{d}\gamma(x,y),$
@@
    Denote conditional independence $X\perp\!\!\!\perp Y~{}|~{}Z$ if $p(x,y|z)=p(x|z)p(y|z)$.
  • Block decomposition. For the parameter $x\in\mathbb{R}^{d}$, we assume there is a fixed block decomposition
    $x=(x_{1},\dots,x_{b}),\quad\forall i\in[b],~{}x_{i}\in\mathbb{R}^{d_{i}},\quad d=\sum_{i=1}^{b}d_{i},$
    where we denote $[b]:=\{1,\dots,b\}$. Denote $-i:=[b]\backslash\{i\}$. For a distribution $\pi$, denote $\pi_{i}$ as the marginal distribution on $x_{i}$, i.e. $\pi_{i}(x_{i})=\int\pi(x_{i},x_{-i})\mathrm{d}x_{-i}$. Denote the partial derivatives $\nabla_{i}=\nabla_{x_{i}},\nabla_{ij}^{2}=\nabla_{x_{i}x_{j}}^{2}$.
2 $\ell_{\infty}$-approximation theory
2.1 $\delta$-locality
We begin by defining the $\delta$-locality of a distribution, which quantifies the locality of the distribution.
@@ 2.1 $\delta$-locality
A distribution $\pi\in\mathcal{P}_{1}(\mathbb{R}^{d})$ is called $\delta$-localized for some constant $\delta>0$, independent of $d$, if for any $i\in[b]$ and $1$-Lipschitz function $\phi:\mathbb{R}^{d_{i}}\rightarrow\mathbb{R}$ such that $\int\phi(x_{i})\pi_{i}(x_{i})\mathrm{d}x_{i}=0$, the solution $u(x)$ to the Stein equation
$\Delta u(x)+\nabla\log\pi(x)\cdot\nabla u(x)=\phi(x_{i})$ (2.1)
@@
2.1 is a technical description of the localized structure of the distribution motivated by the Stein’s method [22]. We will show in 2.4 that the $\delta$-locality condition directly implies that local perturbation of the distribution only introduces dimension independent marginal error. We will see from the proof there that
@@
We keep the technical definition here for the sake of generality. To make it more concrete, we introduce below two common assumptions that imply the $\delta$-locality.
Proposition 2.2.
$\pi\in\mathcal{P}_{1}(\mathbb{R}^{d})$ is $\delta$-localized if it satisfies any of the following conditions:
  1. (1)

### v2

@@ Abstract
Many spatial models exhibit locality structures that effectively reduce their intrinsic dimensionality, enabling efficient approximation and sampling of high-dimensional distributions. However, existing approximation techniques mainly focus on joint distributions, and do not guarantee accuracy for low-dimensional marginals. By leveraging the locality structures, we establish a dimension independent uniform error bound for the marginals of approximate distributions. Inspired by the Stein’s method, we introduce a novel $\delta$-locality condition that quantifies the locality in distributions, and link it to the structural assumptions such as the sparse graphical models. The theoretical guarantee motivates the localization of existing sampling methods, as we illustrate through the localized likelihood-informed subspace method and localized score matching. We show that by leveraging the locality structure, these methods greatly reduce the sample complexity and computational cost via localized and parallel implementations.
@@ 1 Introduction
Approximation and sampling of high-dimensional distributions are fundamental challenges in computational mathematics, statistics and machine learning [27, 18, 23]. A general strategy to mitigate these challenges is to exploit low-dimensional structures in real-world distributions, such as the sparsity [22] and the manifold assumption [16]. In this paper, we study a type of locality structure [4, 35] that models the sparse dependence between components in high-dimensional distributions. This structure arises naturally in spatial models, and has been successfully applied in various fields such as spatial statistics [4, 44, 13], data assimilation [25, 21, 35], and quantum mechanics [28, 3, 6]. Recently, there has been a surging interest for its applications in distribution sampling problems [46, 41, 17, 19, 8]. The goal of this paper is to introduce a quantification of locality structure in high-dimensional distributions, and to develop efficient approximation and sampling methods that exploit this structure to achieve scalability in high dimensions.
Mathematically speaking, the accuracy of various approximation and sampling algorithms are guaranteed by distribution inequalities [34, 20]. For example, denote $1$-Wasserstein distance as $\mathsf{W}_{1}$, use Stein’s Lemma, and one can estabilish an Otto-Villani type inequality [34]
$\mathsf{W}_{1}(\pi,\pi^{\prime})\leq C_{\pi}\mathbb{E}_{\pi^{\prime}}\left[\|\nabla\log\pi-\nabla\log\pi^{\prime}\|\right],$ (1.1)
where $C_{\pi}$ is some constant relates to the log-concavity of $\pi$. Note the right hand side of (1.1) is the difference of the gradient of the log density $\nabla\log\pi$ (the score function), which can be used as a loss function in many distribution estimation and approximation methods, including the score matching and diffusion models [26, 37, 23].
However, (1.1) can be awarkward to use when the distributions are high dimensional. Since the score functions are $d$ dimensional vectors, the right hand side of (1.1) usually grows like $\sqrt{d}$. This may suggest it is necessary to use more training data or finer computational schemes for larger models. The main objective of this paper is to break such pessimistic view by leveraging locality structure. The highlight of the analysis is that the approximation error of low-dimensional marginals can be independent of the ambient dimension, which facilitates efficient approximation and sampling methods [32, 41, 17, 8]. We introduce a $\delta$-locality condition (2.1) to quantify the locality structure, and establish a dimension independent bound for the marginal error of approximate distributions (3.1):
@@
$\max_{i\in[d]}\mathsf{W}_{1}(\pi_{i},\pi_{i}^{\prime})\leq\delta\cdot\max_{j\in[d]}\int\left\|\nabla_{j}\log\pi(x)-\nabla_{j}\log\pi^{\prime}(x)\right\|\pi(x)\mathrm{d}x.$ (1.2)
Here $[d]:=\{1,\dots,d\}$, and $\delta$ is a constant independent of the ambient dimension $d$ for distributions with locality structures, including the sparse graphical models [29].
The inequality (1.2) can be interpretted as a marginal version of (1.1), where the left hand side concerns the marginals of distributions of interest, and the right hand side concerns the difference in the score’s individual components. Using (1.2), we can obtain a dimension independent uniform control of the marginal errors, which is not achievable with (1.1). This is useful if each marginal corresponds to a different location in spatial data [40] and manifold data [42].
Moreover, the marginal form in (1.2) suggests that to obtain a uniform bound on the marginal error, it suffices to minimize the worst coordinate error of the score function, which inspires the localized score matching in Section 5. In particular, we show in Theorem 5.1 the necessary training data can be dimension independent.
The marginal inequality (1.2) is established using Stein’s method [38, 39], which is a useful mathematical tool analyzing the difference between two distributions. To adapt to the marginal setting, we introduce a marginal Stein equation
$\mathcal{L}_{\pi}u(x):=\Delta u(x)+\nabla\log\pi(x)\cdot\nabla u(x)=\phi(x_{i})-\mathbb{E}_{\pi}[\phi(x_{i})],$ (1.3)
where $\phi$ is a test function that only depends on the $i$-th component $x_{i}$. The solution $u$ can then be used to derive the mean difference of $\phi$ on the $x_{i}$-marginals of $\pi$ and another distribution $\pi^{\prime}$. For example, using Stein’s lemma [39], we have
$\mathbb{E}_{\pi^{\prime}}[\phi(x_{i})]-\mathbb{E}_{\pi}[\phi(x_{i})]=\mathbb{E}_{\pi^{\prime}}[\nabla u\cdot(\nabla\log\pi-\nabla\log\pi^{\prime})].$
In Stein’s method, one considers all $\phi$ in a functional class, and seeks a uniform upper bound for all corresponding solution $u$. For example, we will show in 2.1 that if $\pi$ is a sparse graphical model, for all $1$-Lipchitz functions $\phi$, the solution $u$ satisfies $\sum_{j=1}^{d}\left\|\nabla_{j}u\right\|_{L^{\infty}}\leq\delta$, and then we immediately obtain the marginal Otto-Villani inequality
@@
$\mathsf{W}_{1}(\pi_{i},\pi_{i}^{\prime})=\sup_{\left|\phi\right|_{\rm Lip}\leq 1}|\mathbb{E}_{\pi^{\prime}}[\phi(x_{i})]-\mathbb{E}_{\pi}[\phi(x_{i})]|\leq\delta\cdot\max_{j}\mathbb{E}_{\pi^{\prime}}\left[\left\|\nabla_{j}\log\pi-\nabla_{j}\log\pi^{\prime}\right\|\right].$
This motivates the definition of $\delta$-locality, where we use the bound of $u$ to quantify the locality structure. Formally, $u=\mathcal{L}_{\pi}^{-1}(\phi-\mathbb{E}_{\pi}[\phi])$, so that the condition is essentially imposed on $\mathcal{L}_{\pi}^{-1}$, which encodes the locality in $\pi$. A main technical result of this paper is to show that log-concave sparse graphical models are $\delta$-localized with dimension independent $\delta$ (2.1). Thus the $\delta$-locality condition can be regarded a relaxation of the sparse graphical models. We also show that it holds for distributions with certain diagonal dominance property (2.2).
To design efficient approximation and sampling algorithms for localized distributions, a key observation is that the effective dimension of these problems is determined by the sparsity of local dependency rather than the ambient dimension. For instance, by the clique decomposition of the graphical models [4, 29],
$\log\pi(x)=\sum_{j\in[d]}u_{j}(x_{\mathcal{N}_{j}}),$
where $\mathcal{N}_{j}$ is the neighboring indices of $j$. In sparse graphical models, $|\mathcal{N}_{j}|\ll d$, so that the log density is a sum of low-dimensional functions, greatly reducing the complexity of approximating the distribution. Another advantage is that the conditional dependency allow for efficient local and parallel algorithms [41, 17, 19]. In this paper, we discuss two examples of localized sampling methods: the localized likelihood-informed subspace (LIS) method and the localized score matching method, which are exactly the localized versions of LIS methods [11, 10, 12] and score matching [26, 37].
1.1 Related Literature
1.1.1 Locality Structures
Locality structure is an important structure that has been studied accross various fields, including spatial statistics [4, 44, 13], data assimilation [25, 21, 35], and quantum mechanics [28, 3, 6]. In spatial models, it is observed that correlations between parameters diminish significantly at large distances, typically at exponential rate [28, 5, 35]. This allows for computationally efficient approximations by artificially removing long-range dependency.
In spatial statistics, the Vecchia approximation [44] proposes to approximate Gaussian processes by removing conditioners at large distances, i.e.
$\pi(X_{1})\prod_{i=1}^{d-1}\pi(X_{i+1}|X_{1:i})\approx\pi(X_{1})\prod_{i=1}^{d-1}\pi(X_{i+1}|X_{\mathcal{N}_{i+1}}),$
where $\mathcal{N}_{i+1}$ only contains a small subset of preceding points, thus largely reduces the sampling complexity. Various extensions have been proposed based on the nearest-neighbor approximation idea, see the brief review in [13].
In data assimilation, localization methods [25, 21] are introduced to mitigate spurious long-range correlations arising from small ensemble sizes in the ensemble Kalman filter. For instance, the covariance localization artificially remove or dampen long-range correlations in the ensemble covariance $\widehat{C}$, i.e. $\widehat{C}_{\rm loc}=\rho\circ\widehat{C},~{}\rho_{ij}=\psi(|i-j|)$, where $\circ$ denotes the Hadamard product and $\psi:\mathbb{N}\to\mathbb{R}_{+}$ is a rapidly decaying function. Such localization methods have been shown to effectively reduce sampling errors and improve filter accuracy [25, 21, 35].
In quantum mechanics, the nearsightedness principle [28] states the properties of a quantum system are primarily determined by local interactions and are insensitive to distant perturbations. The principle can be formalized as the exponential decay of the density operator $\rho(\bm{r},\bm{r}^{\prime})$ w.r.t. the physical distance $\left\|\bm{r}-\bm{r}^{\prime}\right\|$ [28, 3], which enables efficient sparse approximation of $\rho$. Such approximation can be regarded as the quantum version of localization of distributions, and significantly reduces the computational cost of tasks such as computing the ground states or expectation value of observables in quantum systems. It also lays the theoretical foundation of the interatomic potentials in molecular dynamics simulation [6].
1.1.2 Localized Sampling
In recent years, there has been a fast growing interest in sampling methods that leverage locality structures [46, 32, 31, 41, 19, 17, 8]. [32] proposes to apply the localization technique in Markov chain Monte Carlo (MCMC) to solve Bayesian inverse problems and introduces a localized Metropolis-within-Gibbs sampler. [41] extends this idea and proves a dimension independent convergence rate of the MALA-within-Gibbs for localized distributions, while [31] discusses how to parallelize the proposal computation. Recently, [17] utilizes the localized structure to obtain a dimension independent approximation error for marginals in an image deblurring problem, based on which we develop the idea to the general case in this paper. Meanwhile, [8] demonstrates that when the distribution exhibits localized structure, the bias of the unadjusted Langevin algorithm for the marginals is at most logarithmically dependent on the dimension. Beyond MCMC, [46] proposes Message Passing Stein Variational Gradient Descent (MP-SVGD). By leveraging the conditional independencies, MP-SVGD finds the descent direction coordinate-wisely, and reduces the degeneracy issue of kernel methods in high dimensions. [19] proposes a localized version of the Schrödinger Bridge (SB) sampler. It replaces a single high-dimensional SB problem by $d$ low-dimensional SB problems, avoiding the exponential dependence of the sample complexity on the dimension.
1.1.3 Stein’s Method and Otto-Villani Inequality
Stein’s method is a useful approach for quantifying distances between probability distributions. First developed in [38] for Gaussian approximation, it has been extended to various distributions, including Poisson [7], binomial [39], and high dimensional settings [36, 9]. We refer to Stein’s monograph [39] for a comprehensive review.
Otto and Villani [34] establish a series of inequalities on the relative entropy $\mathsf{H}$, Wasserstein distance $\mathsf{W}_{2}$ and Fisher information $\mathsf{I}$, and reveal their deep connections with the Riemannian-geometric structure of the Wasserstein space of probability measures. The rich structure and wide connections with various fields make itself an important research topic [45]. In this paper, we refer to the Otto-Villani inequality as the “WI” inequality in Definition 3 in [34]:
$\mathsf{W}_{2}(\mu,\nu)\leq C_{\mu}\sqrt{\mathsf{I}(\mu|\nu)},$
where $\mathsf{I}(\mu|\nu)=\mathbb{E}_{\mu}\left\|\nabla\log\frac{\mu}{\nu}\right\|^{2}$ denotes the Fisher information. For more discussions and survey on tranport inequalities, we refer to [45, 20].
1.2 Arrangement and Notations
The paper is organized as follows. We introduce the $\delta$-locality condition in Section 2, and show that it holds for a large class of distributions with locality structure, including the sparse graphical models. In Section 3, we establish the marginal Otto-Villani inequality for $\delta$-localized distributions using the Stein’s method. The next two sections discuss two applications of the theory: localized LIS method in Section 4 and localized score matching in Section 5. In Section 6, we provide the proof of the technical results in Section 2.
@@ 1.2 Arrangement and Notations
  • Sets. Denote $[d]=\{1,\dots,d\}$. Denote the cardinality of a set $A$ as $|A|$.
  • Norms. For $v\in\mathbb{R}^{d}$, denote the vector $p$-norm as $\left\|v\right\|_{p}$ and briefly denote $\left\|v\right\|=\left\|v\right\|_{2}$. Likewisely, for a matrix $A\in\mathbb{R}^{m\times n}$, $\left\|A\right\|=\sup_{v\neq 0}\frac{\left\|Av\right\|}{\left\|v\right\|}$ denotes the matrix $2$-norm. $\left\|A\right\|_{\rm F}=\sqrt{\sum_{i,j}A_{ij}^{2}}$ denotes the Frobenius norm.
    For a scalar function $f:\mathbb{R}^{d}\to\mathbb{R}$, denote $\left\|f\right\|_{L^{p}(\pi)}$ as the weighted $L^{p}$ norm $\left\|f\right\|_{L^{p}(\pi)}=\left(\int_{\mathbb{R}^{d}}|f(x)|^{p}\pi(x)\mathrm{d}x\right)^{1/p}$. For a vector-valued function $f:\mathbb{R}^{d}\to\mathbb{R}^{m}$, similarly denote $\left\|f\right\|_{L^{p}(\pi)}=\left(\int_{\mathbb{R}^{d}}\left\|f(x)\right\|^{p}\pi(x)\mathrm{d}x\right)^{1/p}$. Note here $\left\|f(x)\right\|$ is the vector $2$-norm.
@@
  • Probability. Denote $\mathcal{P}_{1}(\mathbb{R}^{d})$ as the set of probability distributions in $\mathbb{R}^{d}$ with finite first moment. The $1$-Wasserstein distance between two distributions $\pi,\pi^{\prime}\in\mathcal{P}_{1}(\mathbb{R}^{d})$ is
    $\mathsf{W}_{1}(\pi,\pi^{\prime})=\inf_{\gamma\in\Pi(\pi,\pi^{\prime})}\int_{\mathbb{R}^{d}}\left\|x-y\right\|\mathrm{d}\gamma(x,y),$
@@
    Denote conditional independence $X\perp\!\!\!\perp Y~{}|~{}Z\Leftrightarrow p(x,y|z)=p(x|z)p(y|z)$.
2 $\delta$-localized Distributions
In this paper, we consider a high dimensional distribution in $\mathbb{R}^{d}$. In many applications, it is more reasonable to decompose the components into local blocks, and study the dependence between different blocks. Examples include, a small group of genes functioning together as a pathway, and geophysical state of nearby locations representing state of a particular area. To accomodate such settings, we assume there is a fixed block decomposition
$x=(x_{1},\dots,x_{b});\quad\forall i\in[b],~{}x_{i}\in\mathbb{R}^{d_{i}};\quad d=\sum_{i=1}^{b}d_{i},$ (2.1)
where we denote $[b]:=\{1,\dots,b\}$. Denote $-i:=[b]\backslash\{i\}$. For a distribution $\pi$, denote $\pi_{i}$ as the marginal distribution on $x_{i}$, i.e. $\pi_{i}(x_{i})=\int\pi(x_{i},x_{-i})\mathrm{d}x_{-i}$. Denote the partial derivatives $\nabla_{i}f=\nabla_{x_{i}}f\in\mathbb{R}^{d_{i}},\nabla_{ij}^{2}f=\nabla_{x_{i}x_{j}}^{2}f\in\mathbb{R}^{d_{i}\times d_{j}}$.
We begin by defining the $\delta$-locality of a distribution, which quantifies the local parameter dependence of the solution to the marginal Stein equation.
@@ 2 $\delta$-localized Distributions
A distribution $\pi\in\mathcal{P}_{1}(\mathbb{R}^{d})$ is called $\delta$-localized for some constant $\delta>0$, independent of $d$, if for any $i\in[b]$ and $1$-Lipschitz function $\phi:\mathbb{R}^{d_{i}}\rightarrow\mathbb{R}$, the solution $u(x)$ to the marginal Stein equation
$\Delta u(x)+\nabla\log\pi(x)\cdot\nabla u(x)=\phi(x_{i})-\mathbb{E}_{\pi}[\phi(x_{i})]$ (2.2)
@@
2.1 is a technical description of the locality structure of the distribution motivated by the Stein’s method. We will show in 3.1 that the $\delta$-locality condition directly implies that local perturbation of the distribution only introduces dimension independent marginal error. We will see from the proof of 3.1 that
@@
We keep the technical definition here for the sake of generality. In the following, we consider two classes of distributions, i.e. the sparse graphical models and the distributions with certain diagonal dominance condition, and show that they are $\delta$-localized with $\delta$ independent of $d$.
2.1 Sparse Graphical Models
In this section, we show that the sparse graphical model [29], also known as the Markov random field, is $\delta$-localized. As a widely adopted model for distributions with locality structure, it uses an undirected graph $G=(V,E)$ to characterize the local dependence of a random variable $X=(X_{i})_{i\in V}$, where for any nonadjacent vertices $i,j\in V$, the two components $X_{i},X_{j}$ are conditionally independent given the rest of the components $(X_{k})_{k\neq i,j}$.
Precisely, consider an undirected graph $G=(V,E)$ where $|V|=b$. Attach each vertex in $V$ with a unique index $j\in[b]$. Denote $\mathcal{N}_{j}=\{k\in V:(j,k)\in E\}$ as the neighboring vertices of $j$. For simplicity, we require that $E$ includes all the self-loops in $G$, i.e. $(j,j)\in E~{}(\forall j\in G)$.
We say a distribution $\pi(x)\in\mathcal{P}(\mathbb{R}^{d})$ has dependency graph $G$ if for $X\sim\pi$,
$\forall j,k\in V,~{}k\notin\mathcal{N}_{j}~{}\Rightarrow~{}X_{j}\perp\!\!\!\perp X_{k}~{}|~{}X_{\mathcal{N}_{j}^{-}},$ (2.4)
[…]
$\forall k\notin\mathcal{N}_{j}~{}\Rightarrow~{}\nabla_{jk}^{2}\log\pi(x)=0.$
Another equivalent characterization of the sparse graphical model is the clique decomposition [29] of the log density
$\log\pi(x)=\sum_{j\in V}u_{j}(x_{\mathcal{N}_{j}}).$
The clique decomposition is a key property that allows efficient approximation and computation of the distribution, as it implies that the log density is essentially a low-dimensional function, which will be exploited in Section 5.
2.1.1 The Sparsity Condition
To specify the sparsity condition, denote the $q$-th neighbor $\mathcal{N}_{j}^{q}~{}(q\in\mathbb{Z}_{+})$ as
$\mathcal{N}_{j}^{q}:=\{k\in V:\mathsf{d}_{G}(j,k)\leq q\},$
where $\mathsf{d}_{G}(j,k)$ is the graph path distance between $j,k\in V$, i.e. the minimum number of edges that must be traversed to go from $j$ to $k$. For convenience, define $\mathsf{d}_{G}(j,j)=0$, and $\mathsf{d}_{G}(j,k)=\infty$ if $j,k$ are not connected.
Definition 2.2.
An undirected graph $G$ is called $\bm{(S,\nu)}$-local if
$\forall j\in V,~{}q\in\mathbb{N},\quad|\mathcal{N}_{j}^{q}|\leq 1+Sq^{\nu}.$ (2.5)
In the above definition, $S$ denotes the maximal size of the immediate neighbor, and $\nu\in\mathbb{Z}_{+}$ denotes the ambient dimension of the graph, which controls the growth speed of the neighborhood volume with the radius. Here we require it growing at most polynomially to ensure effective locality. A motivating example for 2.2 is the lattice model $\mathbb{Z}^{\nu}$, where a naive bound of the neighborhood volume is
$|\mathcal{N}_{j}^{q}|=|\{k\in\mathbb{Z}^{\nu}:\left\|k\right\|_{1}\leq q\}|\leq(2q+1)^{\nu}<1+(3q)^{\nu}.$
So that $\mathbb{Z}^{\nu}$ is $(3^{\nu},\nu)$-local. A simple class of distributions with such sparse dependency graph is those with banded Hessian, i.e. for some bandwidth $B$, the Hessian $H(x):=\nabla^{2}\log\pi(x)$ satisfies
$\forall|j-k|>B~{}\Rightarrow~{}H_{jk}(x)\equiv 0.$
In this case, the dependency graph is $(2B,1)$-local.
Example 2.1 (1D Ginzburg-Landau).
The Ginzburg-Landau (GL) model is a widely used model in statistical physics [30]. A discrete 1D GL model describes a chain of real-valued spins $\{x_{j}\}_{j=1}^{n}$, where each $x_{j}\in\mathbb{R}$ interacts locally. Its Gibbs distribution is given by
$\pi(x)=\frac{1}{Z}\exp\left(\sum_{j=1}^{n}V(x_{j})+\sum_{j=1}^{n-1}W(x_{j},x_{j+1})\right),$
where $V(x)=\frac{\lambda}{4}(x^{2}-m^{2})^{2}$ is the double-well potential and $W(x,y)=\frac{\beta}{2}(x-y)^{2}$ is the nearest-neighbor interaction. It is clear that $\log\pi$ has banded Hessian with bandwidth $1$, and thus the distribution is $(2,1)$-local.
2.1.2 Sparse Graphical Models are $\delta$-localized
Theorem 2.1.
Let $G=(V,E)$ be a $(S,\nu)$-local graph. Suppose $\pi\in\mathcal{P}(\mathbb{R}^{d})$ has dependency graph $G$, and satisfies for some $0<m\leq M<\infty$,
Then $\pi$ is $\delta$-localized with $\delta=\frac{S\nu!\kappa^{\nu}}{m}$ where $\kappa=\frac{M}{m}$.
Proofs can be found in Section 6.2.
Remark 2.1.
The condition number $\kappa$ plays a crucial role in localization:
  • $\kappa$ is known to be crucial to preserve the band structure in matrix inversion [14, 5]. In probability language, consider a Gaussian distribution $\mathsf{N}(0,C)$, then a moderate $\kappa$ ensures the equivalence of local correlation ($C$ is nearly banded) and conditional dependencies ($C^{-1}$ is nearly banded).
  • To achieve a dimension-independent $\delta$, $\kappa$ must remain bounded as the dimension $d$ grows. This is true in many spatial models where the high dimensionality comes from an extended system rather than finer discretization.
2.2 Diagonal Dominant Distributions
In this section, we consider distributions whose dependency graphs are not necessarily sparse, but instead have a certain diagonal dominant property. Such condition implies that any block is mostly correlated with itself, leading to the localized feature of the distribution. This condition is already studied in [17] motivated by the image deblurring problem, and here we show that it also implies the $\delta$-locality.
Theorem 2.2.